# Dual Target-Mounted RISs-Assisted ISAC Against Eavesdropping and Malicious Interference

### v1

@@ III-B Legitimate RIS Phase Optimization
After relaxing the non-convex constraint (37d) via an SDR-based approach, (P3) can be efficiently solved using the CVX optimization toolbox [29]. However, since the obtained solution is not always guaranteed to be rank-one, additional Gaussian approximation or EVD can be applied to get a feasible rank-one solution [31].
Overall, for $\varepsilon>0$ being the solution accuracy, the complexity of SDR-based solution for beamforming optimization problem (P2) in (34) is calculated as $\mathcal{O}(T_{x}^{4.5}\log(1/\varepsilon)$ while for legitimate phase optimization problem (P3) in (37), it is $\mathcal{O}(N_{L}^{4.5}\log(1/\varepsilon))$ [31].
IV Performance Evaluation
@@
$\theta^{\text{BS}}_{E},\phi^{\text{BS}}_{E}$ $(-160^{\circ},-5^{\circ})$ AoD pairs for $E$-UAV target
@@
Fig. 3 presents the sum secrecy rate of four users, $S_{R}$, as a function of the total transmit power of the BS, $P_{T}$, for $N_{L}=144$ with varying numbers of malicious RIS elements, $N_{M}\in\{49,144,256\}$. These results illustrate the impact of the malicious RIS attacks on $S_{R}$ as $P_{T}$ increases, such that a larger malicious RIS size $N_{M}$ generates stronger interference on the legitimate communication links, and hence $S_{R}$ degrades. Furthermore, in the presence of the malicious RIS (i.e., $N_{M}\in\{49,144,256\}$), it is observed that when $P_{T}$ increases beyond a certain point ($P_{T}=6$ dBW), the $S_{R}$ performance starts degrading. These results can be attributed to the fact that, although a higher $P_{T}$ improves the achievable rate of the legitimate users, it simultaneously increases the sensing interference experienced by the users and enhances the SINR of the $E$-UAV target. This highlights the critical trade-off between enhancing legitimate user rates and mitigating interference effects, which necessitates the appropriate selection of $P_{T}$ values.
@@
Fig. 3 also compares the $S_{R}$ performance of the proposed UAV-mounted RISs-assisted ISAC scheme with that of the benchmark RIS-aided ISAC systems in [24], where one of the targets acts as an eavesdropper without involving a malicious RIS. The results indicate that the proposed system under dual security threats of eavesdropper target and malicious RIS attacks, and the benchmark scheme affected only by an eavesdropper target [24] exhibit a similar $S_{R}$ behavior. As it is expected, the benchmark scheme [24] achieves better $S_{R}$ performance, and performance degradation after a certain point occurs at a higher transmit power ($P_{T}=10$ dBW). This result indicates that the benchmark scheme [24], which operates under a single target eavesdropping security threat, is naturally more robust than the proposed scheme that tackles the more challenging dual-threat scenario of simultaneous eavesdropping and malicious attacks.
@@
In Fig. 4, the $S_{R}$ performance of the proposed scheme with a malicious RIS size of $N_{M}=36$ is evaluated for varying $N_{L}$. The results demonstrate that when $N_{L}$ is significantly larger than $N_{M}$ ($N_{L}=100,121,144$), the $S_{R}$ improves with increasing $P_{T}$, as the abundance of legitimate links offsets the impact of malicious links. However, when the size of $N_{L}$ equals $N_{M}$ ($N_{L}=36$), $S_{R}$ follows an irregular pattern as $P_{T}$ increases, while still maintaining secure communication. This irregularity arises from the comparable strengths of the legitimate and malicious reflections. It can be deduced from the results that $N_{L}$ must be sufficiently larger $N_{M}$ to effectively counteract malicious RIS attacks.
Fig. 5 presents $S_{R}$ performance of the proposed system as a function of $N_{M}\in\{25\sim 225\}$, with $N_{L}=100$ legitimate RIS and varying $P_{T}\in\{-3,0,3,6\}$ dBW. As expected, the results demonstrate that increasing $P_{T}$ improves $S_{R}$ performance. Conversely, a larger $N_{M}$ size intensifies the random interference attacks, which degrade its performance. However, the results also reveal that the proposed scheme sustains secure communication even at relatively low $P_{T}$ values and large $N_{M}$, demonstrating the effectiveness of the proposed algorithm.
@@
In Fig. 6, the $S_{R}$ performance of the proposed scheme for increasing $N_{L}$ sizes is evaluated. It can be clearly observed from the Fig. 6 that an increase in $N_{L}$ enhances the $S_{R}$ performance of the proposed scheme, as it establishes a greater set of legitimate links for communication users. The figure also clearly demonstrates that $S_{R}$ improves as $P_{T}$ increases.
In the following, the sensing performance of the proposed system is evaluated in terms of sensing SINR of both the legitimate and the eavesdropper targets, and the root CRB estimations for AoDs of the eavesdropper UAV target, as illustrated in Figs. 7 and 8, respectively.
In Fig. 7, to evaluate the effectiveness of the proposed algorithm on the sensing SINR performance of UAV targets, the SINR values for both the legitimate and eavesdropper UAVs as a function of $P_{T}$ across varying horizontal distances ($d_{E}^{x}\in 25\sim 50$) of the $E$-UAV are presented. As it can be clear from Fig. 7(a) and 7(b), at a distance of $d_{E}^{x}=25$ m, the UAV targets become at the same distance from the BS (see Table I), yielding similar sensing SINR values for both targets. However, as illustrated in Fig. 7(a), when the $E$-UAV moves farther away, its sensing SINR decreases due to higher path loss, whereas the $L$-UAV’s sensing SINR improves as it experiences less interference from the $E$-UAV.
@@ V Conclusion
In this study, a target-mounted RISs-aided ISAC framework has been proposed to ensure secure communication under dual security threats: an eavesdropper target that intercepts legitimate communications and a malicious RIS that launches random interference attacks. To address these challenges, a non-convex optimization problem is formulated to maximize the sum secrecy rate of the overall system. Then, in order to solve this problem, an SDR-based two-stage algorithm has been developed to optimize the beamforming matrix and phase shifts of the legitimate RIS. Furthermore, comprehensive computer simulations have been conducted to investigate the effectiveness of the proposed algorithm on the secrecy rate and sensing performance metrics across various system configurations. Future research will focus on leveraging RISs to strengthen the security of sensing signals, particularly in dynamic scenarios involving moving targets.

### v2

@@ III-B Legitimate RIS Phase Optimization
After relaxing the non-convex constraint (42d) via an SDR-based approach, (P3) can be efficiently solved using the CVX optimization toolbox [8]. However, since the obtained solution is not always guaranteed to be rank-one, additional Gaussian approximation or EVD can be applied to get a feasible rank-one solution [21]. Overall, after transmit beamforming matrix and legitimate RIS phase shifts are respectively obtained from (38) and (42), the exact achievable secrecy rate of the legitimate system via (31), leveraging the communication and eavesdropper SINR expressions in (28-30).
III-C Malicious RIS Optimization for Worst-Case Scenario
In this section, to evaluate the robustness of the proposed dual target mounted RISs-assisted secure ISAC framework, a worst-case malicious interference scenario is examined in which the hostile system controlling malicious RIS is assumed to have perfect knowledge of the beamforming matrix $\mathbf{W}$, the perfect CSI between the BS–$M$-RIS channel $\mathbf{H}_{M}$ and the $M$-RIS–Uk channel $\mathbf{g}_{M,k}$. Under this worst-case model, the malicious RIS optimizes its phase shift vector to intentionally degrade the received signal power at the legitimate users. Specifically, given the acquired CSI and the legitimate RIS configuration, a dedicated second-stage optimization to minimize the effective channel gain of each legitimate user. This analysis provides insight into the performance limits of the proposed system when subjected to an optimally coordinated interference attack. Then, to disrupt the legitimate communication, the following optimization problem is formulated:
$\text{(P4)}\hskip 8.5359pt\max_{\mathbf{\Omega}_{M}}\hskip 5.69046pt\left\|\mathbf{g}_{M,k}\mathbf{\Omega}_{M}\mathbf{H}_{M}\mathbf{W}\right\|_{\mathrm{F}}^{2}$ (43a)
$\hskip 28.45274pt\text{s.t.}\hskip 5.69046pt|e^{j\zeta_{m}}|=1,\hskip 8.5359pt\text{for}\hskip 5.69046ptm\in\{1,\cdots,N_{M}\}.$ (43b)
Following the optimization of the worst-case malicious RIS interference in (43), its impact on the communication SINR in (28) is incorporated, and the resulting secrecy rate of the legitimate system under these optimized malicious attacks is evaluated using (31).
III-D Computational Complexity
The computational complexity of the proposed scheme is primarily determined by the SDR-based solution employed in the two-stage optimization process. The first stage’s computational complexity of the SDR-based solution for the beamforming optimization problem (P2) in (38) depends on the number of transmit antennas and solution accuracy. For $\varepsilon>0$ being the solution accuracy, the complexity is calculated as $\mathcal{O}(T_{x}^{4.5}\log(1/\varepsilon)$. The second stage’s computational complexity of the SDR-based solution for the legitimate RIS phase optimization (P3) in (42) depends on the number of RIS elements and solution accuracy. Namely, this stage entails a complexity of $\mathcal{O}(N_{L}^{4.5}\log(1/\varepsilon))$ [21].
@@ III-D Computational Complexity
$\tilde{\theta}^{\text{BS}}_{E},\tilde{\phi}^{\text{BS}}_{E}$} $(-160^{\circ},-5^{\circ})$} AoD pairs for $E$-UAV target
$\sigma_{\theta}^{2},\sigma_{\phi}^{2}$} $(5^{\circ})^{2},(5^{\circ})^{2}$} Error variances
@@
IV Performance Evaluation
Figure 3: Beampattern gain of the proposed dual-target RISs-assisted ISAC scheme.
@@ IV Performance Evaluation
Fig. 3 illustrates the beampattern gain of the proposed scheme across both horizontal and vertical angles at $P_{T}=-4$ dBW. Since the beamforming matrix is optimized in (38) to maximize the minimum sensing SINR of the UAV targets, directing the beams primarily toward the $L$-UAV target, it is clearly seen from the results that the beampattern exhibits a distinct peak at the $\theta_{L}^{\mathrm{BS}}$ horizontal angle of $L$-UAV target. Conversely, the gain across the vertical angles remains relatively similar for both targets, as their elevation coordinates are nearly identical.
Fig. 4 illustrates the comparison of the $S_{R}$ performance between the proposed scheme and the RIS-aided ISAC system in [7], which serves as the benchmark. Both schemes adopt the same system configuration with two targets while one target behaves as an eavesdropper. However, no malicious RIS is assumed in benchmark scheme [7]. The results show that the proposed scheme, operating under dual security threats from an eavesdropper target and a malicious RIS attacks, and the benchmark scheme in [7] under a eavesdropper target threat, exhibit similar $S_{R}$ trends, where the benchmark scheme demonstrates a better $S_{R}$ performance. Moreover, it is observed that when $P_{T}$ increases beyond a certain point, the $S_{R}$ performance starts degrading. These results can be attributed to the fact that, although a higher $P_{T}$ improves the achievable rate of the legitimate users, it simultaneously increases the sensing interference experienced by the users and enhances the SINR of the $E$-UAV target. This highlights the critical trade-off between enhancing legitimate user rates and mitigating interference effects, which necessitates the appropriate selection of $P_{T}$ values. Furthermore, it can be deduced from the results that the proposed scheme sustains reliable communication under increasing malicious attacks.
@@
Fig. 5 illustrates the $S_{R}$ performance of the proposed scheme with a fixed malicious RIS size of $N_{M}=36$, which is evaluated for varying $N_{L}$. As expected, the secrecy rate generally improves with higher $P_{T}$, which provides a larger power budget to simultaneously meet sensing constraints and overcome interference. Moreover, increasing $N_{L}$ yields substantial performance gains. A larger legitimate RIS provides greater spatial degrees of freedom and higher passive beamforming resolution.
In Fig. 6, $S_{R}$ performance of the proposed system for $N_{L}=100$ and different $N_{M}\in\{25\sim 225\}$ is presented. Although the improvement is modest at low $P_{T}$, the results show that $S_{R}$ of the proposed system improves as $P_{T}$ increases. Moreover, it can be observed that larger values of $N_{M}$ lead to a degradation in the secrecy rate performance. The results also reveal that the proposed scheme maintains secure communication even at relatively low $P_{T}$ values and large $N_{M}$ cases, thereby demonstrating the robustness of the proposed algorithm against both eavesdropping and malicious attacks.
@@
Fig. 7 illustrates the $S_{R}$ performance of the proposed scheme for increasing $N_{L}$ sizes. As observed, the secrecy rate increases with $N_{L}$ for all considered power levels. This highlights that scaling the legitimate RIS provides spatial degrees of freedom, translating to a higher passive beamforming gain. Consequently, the legitimate system can more effectively direct signals to intended users, even when the transmit power is strictly limited (e.g., at $P_{T}=-6$ dBW). Furthermore, an increase in $P_{T}$ improves the system’s performance across all $N_{L}$ configurations, as the BS gains a larger power margin to simultaneously satisfy the dual objectives of secure communication and high-resolution sensing.
In the following, the sensing performance of the proposed system is evaluated in terms of sensing SINR of both the legitimate and the eavesdropper targets, and the root CRB estimations for AoDs of the eavesdropper UAV target, as illustrated in Figs. 8 and 9, respectively. It should be emphasized that the CRB is utilized in the results as a conclusive evaluation benchmark to validate the effectiveness of the SINR-based beamforming design, rather than serving as an active constraint during the optimization phase.
In Fig. 8, to evaluate the effectiveness of the proposed algorithm on the sensing SINR performance of UAV targets, the SINR values for both the legitimate and eavesdropper UAVs as a function of $P_{T}$ across varying horizontal distances ($d_{E}^{x}\in 25\sim 55$) of the $E$-UAV are presented. As it can be clear from Fig. 8(a) and 8(b), at a distance of $d_{E}^{x}=25$ m, the UAV targets become at the same distance from the BS (see Table I), yielding similar sensing SINR values for both targets. However, as illustrated in Fig. 8(a), when the $E$-UAV moves farther away, its sensing SINR decreases due to higher path loss, whereas the $L$-UAV’s sensing SINR improves as it experiences less interference from the $E$-UAV.
@@
Figure 10: Secrecy rate performance of the worst-case malicious interference.
In Fig. 10, $S_{R}$ performance of the proposed scheme is illustrated for both random malicious interference and optimized interference under the worst-case scenario, where the transmit beamforming and the CSI between the BS and the $E$-UAV are compromised and perfectly known by the $E$-UAV. It is clear from the results that increasing $N_{L}$ improves the $S_{R}$ performance for both cases. Moreover, the worst-case malicious RIS interference yields a significantly lower $S_{R}$ than random interference, demonstrating the effectiveness of our optimization algorithm in (43). Specifically, the optimized interference strategy exploits the compromised beamforming and CSI knowledge to intentionally degrade the $S_{R}$, thereby representing a stringent benchmark scenario. Nevertheless, the results indicate that a positive $S_{R}$ is still achievable even under worst-case conditions, confirming that reliable and secure communication remains feasible despite strong adversarial interference.
@@ V Conclusion
In this study, a dual target-mounted RISs-aided ISAC framework has been proposed to ensure secure communication under dual security threats: an eavesdropper target that intercepts legitimate communications and a malicious RIS that launches random interference attacks. To address these challenges, a non-convex optimization problem is formulated to maximize the sum secrecy rate of the overall system. Then, in order to solve this problem, an SDR-based two-stage algorithm has been developed to optimize the beamforming matrix and phase shifts of the legitimate RIS. Moreover, a worst-case malicious interference scenario has been investigated, in which the malicious RIS interference has been optimized to maximally degrade the secrecy rate of the legitimate system, thereby evaluating the robustness of the proposed framework under the most adversarial conditions. Furthermore, comprehensive computer simulations have been conducted to investigate the effectiveness of the proposed algorithms on the secrecy rate and sensing performance metrics across various system configurations. Future research will focus on leveraging RISs to strengthen the security of sensing signals, particularly in dynamic scenarios involving moving targets.